OpenStudy (anonymous):

Trig find the exact value of the expressions cos(a+B), sin(a+B) and tan(a+B) under the following conditions: sin(a)=(40/41), a lies in quadrant I, and sin(B)=(24/25), B lies in quadrant II. 1. cos(a+B)= 2. sin(a+B)= 3. tan(a+B)=

Nnesha (nnesha):

40/41 or 40 pi/41 ?

OpenStudy (anonymous):

40/41

OpenStudy (anonymous):

¯\_(ツ)_/¯

OpenStudy (isaiah.feynman):

You need to use trig identities.

OpenStudy (isaiah.feynman):

\[\sin(a+b) = \sin(a)\cos(b)+\sin(b)\cos(a)\]

OpenStudy (anonymous):